solve the equation by completing the square. the equations has real number solutions.
3x^2-2x-56=0

Answer by lwsshak3(11628)

solve the equation by completing the square. the equations has real number solutions.
3x^2-2x-56=0
...
3x^2-2x-56=0
completing the square
3(x^2-2x/3+1/9)-56-1/3=0
3(x-1/3)^2=56+1/3=168/3+1/3=169/3
divide by 3
(x-1/3)^2=169/9
take sqrt of both sides
x-1/3=±√169/3
x=1/3±√169/3
x=(1+√169)/3=(1+13)/3=14/3
or
x=(1-√169)/3=(1-13)/3=-12/3=-4
Check:
for x=-4
3x^2-2x-56=0
3(4^2)-2(-4)-56=48+8-56=0
..
for x=14/3
3(14/3)^-2(14/3)-56=65.33-9.33-56=0